A project has an initial cost of $43,175 expected net cash inflows of 11,000 per year for 10 years and a cost of capital of 10% what is the projects PI
Answer #1

Initial Investment = C = $ 43175 and Net Annual Cash Inflows = $ 11000, Cost of Capital = 10 % and Tenure = 10 Years

Total Present Value of Net Annual Cash Inflows = B = 11000 x (1/0.1) x [1-{1/(1.1)^(10)}} = $ 67590.24

Profitability Index = PI = B / C = 67590.24 / 43175 ~ 1.56549 ~ 1.56